40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 2 18:35:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
    Reply  Tue Feb 2 18:54:02 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
       Reply  Wed Feb 3 23:23:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table KnifeEdgeScans_x.pdfBeamscan_x.pdf
          Reply  Thu Feb 4 09:16:46 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
             Reply  Fri Feb 5 00:23:01 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table Beamscan_setup.pdfBeamscan_x.pdfBeamscan_y.pdfZscan.pdf
                Reply  Fri Feb 5 16:50:24 2016, gautam, Update, Green Locking, First pass at mode-matching Modematch_AUXx.pdfNewSetUp.pngModematch_AUXx_2.pdfXendModeMatch.m.zip
                   Reply  Mon Feb 8 15:36:37 2016, gautam, Update, Green Locking, Alternative mode-matching scheme Modematch_alternative.pdf
                      Reply  Tue Feb 9 04:37:10 2016, ericq, Update, Green Locking, Laser swap initiated 
                         Reply  Tue Feb 9 11:49:47 2016, gautam, Update, Green Locking, Laser swap initiated Innolight_beamProfile.pdf
                            Reply  Tue Feb 9 19:15:36 2016, gautam, Update, Green Locking, Laser swap - updates Modematch_X.pdfXendModeMatch.m.zip
                               Reply  Wed Feb 10 17:57:15 2016, gautam, Update, Green Locking, Laser swap - updates 
                                  Reply  Fri Feb 12 17:05:40 2016, gautam, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                     Reply  Fri Feb 12 19:07:52 2016, Koji, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                        Reply  Fri Feb 19 00:42:19 2016, gautam, Update, Green Locking, Laser swap - beam ellipticity from laser? IMG_6287.JPGIMG_6288.JPGIMG_6286.JPG
                                           Reply  Mon Feb 22 13:56:52 2016, gautam, Update, Green Locking, Laser swap -reflected beam from ETM aligned 
                                              Reply  Tue Feb 23 23:01:16 2016, gautam, Update, Green Locking, Laser swap - Green PDH locking AM_scan.pdfAM_scan_zoomed.pdf
                                                 Reply  Wed Feb 24 19:29:13 2016, gautam, Update, Green Locking, Laser swap - Green PDH locked 
                                                    Reply  Fri Feb 26 01:52:44 2016, gautam, Update, Green Locking, Laser swap - Green PDH OLTF X_PDH_OLTF_20160225.pdf
                                                       Reply  Mon Feb 29 17:17:26 2016, gautam, Update, Green Locking, Laser swap - still no green beatnote 
                                                          Reply  Wed Mar 2 17:42:19 2016, gautam, Update, Green Locking, Laser swap - some progress 
                                                             Reply  Thu Mar 3 01:25:50 2016, gautam, Update, Green Locking, Laser swap - 2 IR + 1 green beatnotes found IR_beat_20160303_2.pdf
                                                                Reply  Fri Mar 4 01:11:41 2016, gautam, Update, Green Locking, Laser swap - both green beatnotes found IR_beat_20160303_green.pdf
                                                                   Reply  Sat Mar 5 23:31:01 2016, gautam, Update, Green Locking, Laser swap - some updates AUXxTelescope.png.pngPM_setup.pdfPLLolg.pdfAMPM20160303.pdfIRbeat_20160304.pdf
                                                                      Reply  Mon Mar 7 23:51:36 2016, gautam, Update, Green Locking, Laser swap - some improvement PomonaTF.pdfXPDH.pdfYPDH.pdfgreenbeat_20160307.pdf
                                                                         Reply  Tue Mar 8 18:22:20 2016, rana, Update, Green Locking, Laser swap - some improvement 
Message ID: 12026     Entry time: Mon Mar 7 23:51:36 2016     In reply to: 12023     Reply to this: 12027
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Laser swap - some improvement 
Quote:

 

Next steps in recovering ALS and trying to lock again

  • Having set the PDH modulation frequency to 256.62kHz, I took the spectrum of ALS noise using the IR beat (i.e. by piping the IR beat signal through the electronics the green beats usually go through - 6dB and 10dB attenuators were placed immediately after the beat PDs for the X and Y arms respectively, to make the signal levels compatible with the electronics), Attachment #5 unfortunately suggests that the noise performance is still poor, and I suspect the situation will be similar using the green beat (though I have not measured this yet).
  • The modulation depth could be sub-optimal for the X-end PDH, I have to measure this and check that it is at an acceptable level. This will also tell me if I need to change the sum+HPF pomona box used to send the PDH control signal + piezo dither signal to the laser PZT. In order to do this, I need to know what the input impedance to the FAST control BNC is - the manual isn't very helpful, it just says the piezo has a capacitance less than 10,000pF. I suppose I will have to actually measure this.
  • PDH loop OLTFs have to be re-measured for both ends to check that the servo gain's are appropriately placed.
  • We know that the mode-matching into the arm for the X end is poor (I have yet to quantify this) - I suspect that the beam ellipticity is the main culprit. However, the DC transmitted power levels at the PSL table are comparable to (even slightly better than) the Y arm numbers, and so this cannot be the sole reason why the X-arm ALS noise is so much worse... I will continue my investigations next week...

Attachment #1

Since I could not determine how many volts at the LO input of the pomona box input corresponds to how many volts at the laser PZT, I measured the transfer function between these points using the Agilent network analyzer. The measured TF suggests that for a function generator output of 2Vpp, we get approximately 75mrad of phase modulation, which compares reasonably well with the value of 120mrad reported here. I did not attempt to further increase the LO output signal to push this number closer to 120mrad, as with 2Vpp from the function generator we get +7dBm at the mixer, which is what it wants - so I wanted to avoid any attenuators etc...

Attachments #2 and #3

After ensuring that we have appreciable phase modulation, I set out to measure the PDH OLTFs and adjust the gain on the uPDH boxes accordingly. The X end gain is at 6.0, and the Y end gain is at 4.0. Before measuring the Y-end OLTF, I adjusted the steering mirrors to increase GTRY to ~0.45. GTRX remains a paltry 0.05... But the UGFs seem satisfactory..

Attachment #4

Finally, I took the ALS noise spectrum for the green beats. The beat note amplitudes on the network analyzer in the control room are still puny compared to what we had, -40dBm for Y and -45dBm for X. But the phase tracker Q values are ~1000 and ~3000 for X and Y respectively, which are pretty close to what these were if memory serves me right. There may still be some room for optimization of the PDH loop gains etc, and we could perhaps look at lowering the gain of the REFL PD at the X end? I also have yet to do the sweep for the 3 temperatures at which we can find a beatnote and park at the middle one...

These spectra suggest we could even possibly try locking? We are approximately a factor of 3 above the reference for X and on par with the reference for Y....

Unrelated to this work: I also realinged the PMC, PMC transmission is now 0.730V up from ~0.65V.

Attachment 1: PomonaTF.pdf  55 kB  | Hide | Hide all
PomonaTF.pdf
Attachment 2: XPDH.pdf  52 kB  | Hide | Hide all
XPDH.pdf
Attachment 3: YPDH.pdf  53 kB  | Hide | Hide all
YPDH.pdf
Attachment 4: greenbeat_20160307.pdf  62 kB  | Hide | Hide all
greenbeat_20160307.pdf
ELOG V3.1.3-